OpenStudy (anonymous):

[Statistics] If n=5, x=0, p=.3, then p(x) = (n/x)(p)^(x)(q)^(n-x). Help :(

OpenStudy (shamil98):

that looks scary.

OpenStudy (shamil98):

Is it just evaluating?..

OpenStudy (anonymous):

Yes :'(

OpenStudy (anonymous):

If n=5, x=0, p=.3, then p(x) = (n/x)(p)^(x)(q)^(n-x). Do you want to find the value at x= 0 @ifunfrank

OpenStudy (anonymous):

Yes, but I don't think it lets me go through when x is in the denominator

OpenStudy (anonymous):

Yes because n/x= 5/ 0 = becomes indeterminate i.e. \[\frac {n}{x} = \infty \] and the answer... @ifunfrank

OpenStudy (anonymous):

\[p(x) = (\frac{ n }{ x })(p)^x(q)^{n-x}\] Here is a clearer format of the equation

OpenStudy (anonymous):

@ifunfrank Again this is the same as given above...it will also give the same answer. since x is again as a denominator.

OpenStudy (anonymous):

I need an answer to the fourth decimal place and when I put infinity as the answer it says it's incorrect

OpenStudy (anonymous):

most probably question could be wrong.